OpenStudy (anonymous):

If x:6 as 3:9, then x is equal to____ a0.? please help meee!!!!

OpenStudy (anonymous):

2, because 3=9/3 and x=6/3 which is 2.

OpenStudy (anonymous):

thanks!!!!!!!!!

OpenStudy (anonymous):

Your welcome :)